# On the relevance of pre-neural approaches in natural language processing pedagogy

### v1

@@ Abstract
While neural approaches using deep learning are the state-of-the-art for natural language processing (NLP) today, pre-neural algorithms and approaches still find a place in NLP textbooks and courses of recent years. In this paper, we compare two introductory NLP courses taught in Australia and India, and examine how Transformer and pre-neural approaches are balanced within the lecture plan and assessments of the courses. We also draw parallels with the objects-first and objects-later debate in CS1 education. We observe that pre-neural approaches add value to student learning by building an intuitive understanding of NLP problems, potential solutions and even Transformer-based models themselves. Despite pre-neural approaches not being state-of-the-art, the paper makes a case for their inclusion in NLP courses today.
@@ 1 Introduction
Transformer-based models are the state-of-the-art in natural language processing (NLP). They represent a new era of neural approaches [^1] The phrase ‘neural approaches’ is expected to refer to the broad spectrum of neural approaches using deep learning. This includes dense word representations, Transformer and Transformer-based models used in NLP. for NLP. From early work in word representations to recent approaches in instruction tuning, neural approaches have significantly transformed NLP and many other areas of artificial intelligence. Universities around the world have incorporated neural approaches in their introductory NLP courses, arguably with varying speed, and to different extents. This puts the relevance of pre-neural approaches in NLP courses of today under question. This reflection from academics/faculty members may also be based on questions from students taking the course who wonder why they should learn about pre-neural approaches.
Therefore, this paper analyses the role of pre-neural approaches in the context of modern university-based courses in NLP. We focus on introductory NLP courses offered to students of computer science & engineering or equivalent degrees, and investigate the question:
How can an introductory NLP course balance between the content covering neural and pre-neural approaches?"
We address the question in two parts: (a) how others do it, and (b) how we do it. With respect to (a), we describe summaries of NLP textbooks and publicly available courses in the context of the question. As for (b), we draw insights from our experience teaching NLP courses at two large, research-intensive universities in Australia and India. We introduce and discuss our motivations, considerations and decisions in the lectures, tutorials and projects of NLP courses. Authors of this paper represent two personas of NLP educators: (a) an early-career educator who introduced a new NLP course in the beginning of 2024; and (b) a seasoned educator with two-decade of teaching and research experience. Therefore, the main focus of the paper is to compare the NLP courses run at the two universities.
The novelty of this paper is as follows:
@@
    There have been papers in the past on specific aspects (tutorials, lecture content, etc.) of individual NLP courses Plank (2021); Foster and Wagner (2021); Gaddy et al. (2021). This paper is novel in its comparison of two NLP courses taught by educators with significantly different experience in NLP.
@@
    We also draw parallels to the age-long objects-first or -later debate in computer science education, to highlight that the (pre-)neural discussion in NLP courses can benefit from lessons in computer science education.
@@
    We hope that the paper serves as a useful resource for NLP educators to examine the role of pre-neural methods in their curricula now and in the future.
The rest of the paper is described as follows. We first introduce the context of the paper in terms of the two courses being compared in Section 2. Following that, we discuss an analogous objects first- and objects-later- debate in computer science education in Section 3. We describe how NLP textbooks and NLP courses (as per publicly available course outlines) cover pre-neural approaches in Section 4. We then proceed to compare the lecture plan, and coding assessments in the two courses in Sections 5 and  6 respectively. Section 7 puts it all together to make our case for the relevance of pre-neural approaches in NLP courses. Finally, Section 8 concludes the paper.
@@ 2 Context
This paper is based on experiences from NLP courses run at two large universities in Australia and India, both titled ‘Natural Language Processing’. We refer to these as NLP-UNSW and NLP-IITB. NLP-UNSW is the first offering of the course with course strength of 60, while NLP-IITB is the ninteenth offering of the course with course strength of 150. NLP-IITB is accompanied with a ‘Deep learning for NLP’ course which often follows NLP-IITB. Both courses were offered as an undergraduate/postgraduate elective, and were an introductory course to NLP-IITB. The cohorts consisted of undergraduate (nearly 50%) and postgraduate students. Postgraduate students included those enrolled in Masters by Research programme and early PhD students.
The course components of NLP-UNSW and NLP-IITB are in Figures 1 and  2 respectively. The text in red highlights the course convener’s reasons behind covering pre-neural approaches, and shows a significant overlap. Generally speaking, it allowed the students to solve problems and understand deep learning approaches vis-a-vis their predecessors. NLP-UNSW insisted that the group project compare deep learning methods with their predecessors, NLP-IITB allowed the students to choose the right method. In contrast, the individual assignment in NLP-UNSW only involved using NLP libraries while NLP-IITB used multiple assignments covering both statistical as well as neural models.
@@ 3 Parallels to Computer Science Education
As NLP educators grapple with Transformers-first, -later, or perhaps -interleaved approaches to curricula, it is helpful to draw upon empirical research and lessons learned by CS1 educators (introductory programming) who have grappled with a similar dilemma: the age-long objects-first or -later debate. For decades, introductory programming students learned imperative-style programming in languages such as C, Pascal, and Fortran. In the early 2000s, educators started incorporated object-oriented programming concepts into their curriculum as object-oriented programming took hold in industry Cooper et al. (2003). Where the objects-first community may begin immediately with concepts of composition, inheritance, and abstraction, the objects-later community instead focus on simpler, foundational programs and concepts such as sequence and control flow.
@@
Unlike CS1 education, however, NLP may often be available as a single elective in a computer science programme. Therefore, NLP educators are forced to balance providing practical, post-transformer skills to students while also providing solid foundations in NLP in a shorter time-frame, perhaps supporting an interleaved approach to NLP curricula.
Authors; Publisher Year Layout Deep learning focus
Speech and Language Processing; Dan Jurafsky, James H Martin; - 2024 Fundamental Algorithms (Statistical Models, Neural Models, RNNs, LSTMs.) followed by NLP applications and linguistic tasks. Techniques first. Pre-neural: Yes.
@@
Natural Language Processing; Pushpak Bhattacharyya, Aditya Joshi; Wiley 2023 Fundamental techniques first. Then neural. Significant focus on NLP problems. Interleave pre-neural and neural both. Three generations.
@@
Practical Natural Language Processing; Sowmya Vajjala, Bodhisattwa Majumder, Anuj Gupta, Harshit Surana; O’Reilly 2020 NLP Primers: Quick introduction to primary NLP concepts. Vey task driven: Classification, IE, Chatbots, Applications to domains Cover deep learning Early. Focus on applications and tasks
@@
Table 1: Deep learning focus in NLP textbooks listed in the reverse-chronological order.
@@ 4 Observations from NLP textbooks and courses
Table 1 presents a summary of some NLP textbooks, in terms of layouts and their deep learning focus. We observe that nearly all textbooks cover pre-neural approaches, primarily in terms of statistical models. We describe two textbooks in particular. The textbook by Jurafsky-Martin  [^2] https://web.stanford.edu/~jurafsky/slp3/ divides its chapters into fundamental algorithms and NLP applications. The first five chapters introduce regular expressions, one-hot vectors and statistical algorithms like support vector machines and logistic regression. In the subsequent chapters, the book covers word embeddings, then begins to combine LSTM with CRFs in the context of POS tagging. Following that, the book covers Transformer, fine-tuning and prompting/prompt learning Transformer. Bhattacharyya and Joshi (2023) use a different approach. They cover fundamental algorithms for representation learning: computational grammar, probabilistic language modeling and word2vec/LSTM-based representations. The book visualizes NLP as three generations: rule-based, statistical and neural. Following that, the book introduces Transformer. The subsequent chapters alternate between the approaches in the three generations.
Table 2 illustrates how some NLP courses cover pre-neural topics. We only use courses whose course outlines are publicly available on the internet. Similarly, this is not a complete list either. In universities where there is a deep learning-focussed course for NLP, topics such as syntactic parsing or statistical classification are covered in substantial detail. Even for courses that cover little pre-neural content, n-gram language modeling is a topic that is included. The table also shows ‘when is Transformer taught’ which may be viewed as a turning point when deep learning-based approaches assume focus in the course.
@@ 5 Lecture Plan
In the NLP-UNSW course which is spread over 10 weeks, we adopt a hybrid methodology where we interleave neural approaches with statistical/rule-based approaches. The hybrid methodology is illustrated in Figure 3. In the first week, we introduce NLP via black-boxes such as spacy, nltk and HuggingFace pipelines. The students are introduced to NLP tasks with demonstrations of the three libraries. This is to help the students develop an understanding of the task. Similarly, if students aim to only ‘use’ NLP in their projects (which may be in other disciplines), these libraries are sufficient. The focus here is also to help them understand the input and output of NLP systems, which we believe is the starting point to understand NLP.
@@
In week 2, we cover representations of text via one-hot vectors and probabilistic language modeling (statistical approach) and word representation learning as in word2vec and GloVe (neural approach). This not only allows students to appreciate the value addition of neural approaches but also identify situations in which non-neural approaches may be sufficient. In week 3, we focus on Transformer architecture: exposing the students to the architectural details, pseudocode and code implementations. This is followed by Transformer-based models (encoder, decoder and encoder-decoder models) in week 4. With this background in neural NLP models, we switch gears and focus on one NLP task every week. Every task is covered using the following steps: the linguistic task and associated challenges, pre-neural approaches, neural approaches, and recent advances in the area. The first two allow the students to gain a foundational understanding of the problem, followed by the state-of-the-art in deep learning. The recent advances are catered to students who might be interested in future research in NLP without going into too much depth - to ensure that the content is accessible. The NLP tasks we cover are: sentiment analysis (representing sequence classification), named entity recognition (representing token classification), machine translation (representing sequence-to-sequence tasks), summarization (modified sequence-to-sequence tasks) and bias mitigation.
@@
NLP-IITB starts with focus on sequence labeling via HMM-based POS tagging, followed by tree extraction via probabilistic parsing. In both the cases, the mathematical details (driven by probability) go hand in hand with the algorithms (explained through the code). Then, machine translation is introduced, which makes way for an introduction to Transformer and large language models. The subsequent weeks discuss sentiment analysis as a specific NLP task, and evaluation metrics in NLP. It must be noted that NLP-IITB is the introductory NLP course while there is a deep learning-specific NLP course at IITB which is often a follow-up to the introductory NLP-IITB. This is not the case for NLP-UNSW (being the only NLP-centric course). As a result, the course needs to balance between the two. This comparison shows how educators may make different choices, depending on the length of the teaching period. As a result, the lecture plan of NLP-IITB uses the spectrum shown in Figure 4. Pre-neural approaches allow focusing on the linguistic phenomenon which is captured using a model. The data provides the parameters of the model.
@@
In both NLP-UNSW and NLP-IITB, however, we observe that the foundational material serves as a good basis to introduce terminology (for example, types of summarization are abstractive and extractive summarization) and tag sets (as in the case of named entity recognition). The pre-neural approaches build an intuitive understanding of building computational models for the specific NLP task. Appreciating the challenges in the pre-neural approaches lead the students to the need for deep learning approaches. Through handwritten examples in the lecture, we highlight the relationships between the mathematics of pre-neural approaches and Transformer-based approaches (for example, alignment model in statistical MT with cross-attention in Transformer).
@@ 6 Coding Assessments
Both NLP-UNSW and NLP-IITB have individual and group coding assessments. NLP-UNSW has one of each, while NLP-IITB has two individual and one group assessment.
@@ 6.1 Course Assignment
The first coding assessment in NLP-UNSW is an individual coding assignment completed by students by the end of week 3. The coding assessment is run in a competition-like environment. The students are given a document defining the problem definition, and a template code that also contains stubs for testing. The students are required to add to the template. Closer to the submission, the students are given test cases to evaluate their code. The topic of the individual assignment needed deliberation. Until that point, we have only started discussing Transformer in the lectures. Therefore, we utilize the individual assignment to assess the pre-neural skills of the students. In the case of this offering, the task was to extract skills from job ads, based on a skills ontology. The assignment was designed in two parts: the first part required the students to use the spacy matcher, while the second part required them to use embedding similarities. The former tests them for their ability to use a pre-neural library while the latter requires them to use HuggingFace models. As a result, the assessment covers the students’ ability to ‘use’ NLP models. Marking the assignment includes an automatic marking component with test cases and manual marking by the tutors. Being a rule-based system, the evaluation of test cases is found to return low precision. Therefore, rather than scoring on individual test cases, we mark the students on precision thresholds.
NLP-IITB follows a similar strategy although there are multiple individual assignments. The course convener found it useful to give pre-neural assignments on topics such as POS tagging, parsing and so on. An innovative assignment was also tried out, which met with good student feedback. Given a large dataset of names of cities, the task was to identify a preferred suffix in the name of a city. Students reported multiple deep learning/non-deep learning approaches.

### v2

# Striking a Balance between Classical and Deep Learning Approaches in Natural Language Processing Pedagogy

@@ Abstract
While deep learning approaches represent the state-of-the-art of natural language processing (NLP) today, classical algorithms and approaches still find a place in NLP textbooks and courses of recent years. This paper discusses the perspectives of conveners of two introductory NLP courses taught in Australia and India, and examines how classical and deep learning approaches can be balanced within the lecture plan and assessments of the courses. We also draw parallels with the objects-first and objects-later debate in CS1 education. We observe that teaching classical approaches adds value to student learning by building an intuitive understanding of NLP problems, potential solutions, and even deep learning models themselves. Despite classical approaches not being state-of-the-art, the paper makes a case for their inclusion in NLP courses today.
@@ 1 Introduction
Transformer-based models are the state-of-the-art in natural language processing (NLP). They represent a new era of deep learning approaches [^1] The phrase ‘deep learning approaches’ is expected to refer to the broad spectrum including but not limited to dense word representations, Transformer and Transformer-based models used in NLP. for NLP. From early work in word representations to recent approaches in instruction tuning, deep learning approaches have significantly transformed NLP and many other areas of artificial intelligence. Universities around the world have incorporated deep learning approaches in their introductory NLP courses, arguably with varying speed, and to different extents, as evidenced to an extent, in Table 2). This puts under question the relevance of classical approaches, i.e., those pre-dating deep learning approaches, in NLP curricula. This reflection from academics/faculty members may also be based on questions from students taking the course who wonder why they should learn about classical approaches, as seen in pre-course feedback received by the authors of this paper.
Therefore, we analyse the role of classical approaches in the context of modern university-based courses in NLP. We focus on introductory NLP courses offered to students of computer science & engineering or equivalent degrees, and investigate the question:
How can an introductory NLP course balance between the content covering classical and deep learning approaches?"
We address the question in two parts: (a) how others do it, and (b) how we do it. With respect to (a), we describe summaries of NLP textbooks and publicly available courses in the context of the question. As for (b), we draw insights from our experience teaching NLP courses at two large, research-intensive universities in Australia and India. We introduce and discuss our motivations, considerations and decisions in the lectures, tutorials and projects of NLP courses. Authors of this paper represent two personas of NLP educators: (a) an early-career educator who introduced a new NLP course in the beginning of 2024; and (b) a seasoned educator with two-decade experience of teaching and research experience. The novelty of this paper is as follows:
@@
    There have been papers in the past on specific aspects (tutorials, lecture content, etc.) of individual NLP courses Plank (2021); Foster and Wagner (2021); Gaddy et al. (2021). This paper is novel in its comparison of two NLP courses taught by educators with significantly different experience in NLP research and pedagogy.
@@
    We also draw parallels to the age-long objects-first or -later debate in computer science education, to highlight that the classical approaches in NLP courses can benefit from lessons in computer science education.
@@
    We hope that the paper serves as a useful resource for NLP educators to examine the role of classical methods in their curricula now and in the future.
The rest of the paper is described as follows. We first introduce the context in terms of the two courses being compared in Section 2. Following that, we discuss an analogous objects first- and objects-later- debate in computer science education in Section 3. We describe how NLP textbooks and NLP courses (as per publicly available course outlines) cover classical approaches in Section 4. We then proceed to compare the lecture plan, and coding assessments involved in the two courses in Sections 5 and  6 respectively. Section 7 puts it all together to make our case for the relevance of classical approaches in NLP courses. Finally, Section 8 concludes the paper.
@@ 2 Context
The two NLP courses we compare are run at two large universities in Australia and India, both titled ‘Natural Language Processing’. We refer to these as NLP-UNSW and NLP-IITB [^2] UNSW: University of New South Wales, Sydney, Australia; IITB: Indian Institute of Technology Bombay, Mumbai, India.. NLP-UNSW is the first offering of the course with 60 enrolled students, while NLP-IITB is the ninteenth offering of the course with 150 enrolled students. While NLP-UNSW is the only NLP-focussed course at UNSW at the time of writing this paper, NLP-IITB is accompanied with a ‘Deep learning for NLP’ course which often follows NLP-IITB. Both courses were offered as an undergraduate/postgraduate elective as a part of the computer science and engineering programme, and were introductory courses to NLP. The cohorts consisted of undergraduate (nearly 50%) and postgraduate students. Postgraduate students included those enrolled in Masters by Coursework or Masters by Research programmes, and early PhD students.
The course components of NLP-UNSW and NLP-IITB, along with the course conveners’ reasons behind covering classical approaches, are shown in Figures 1 and  2 respectively. Generally speaking, an exposition to classical approaches allowed the students to solve problems and understand deep learning approaches vis-á-vis their predecessors. NLP-UNSW insisted that the group project compare deep learning methods with their predecessors. NLP-IITB allowed the students to choose an appropriate method. In contrast, the individual assignment in NLP-UNSW only involved using black-box NLP libraries while NLP-IITB used multiple assignments covering both statistical as well as deep learning models.
@@ 3 Parallels to Computer Science Education
As NLP educators grapple with Classical-first, -later, or perhaps -interleaved approaches to curricula, it is helpful to draw upon empirical research and lessons learned by CS1 educators (introductory programming) who have grappled with a similar dilemma: the age-long objects-first or -later debate. For decades, introductory programming students learned imperative-style programming in languages such as C, Pascal, and Fortran. In the early 2000s, educators started incorporated object-oriented programming concepts into their curriculum as object-oriented programming took hold in industry Cooper et al. (2003). Where the objects-first community may begin immediately with concepts of composition, inheritance, and abstraction, the objects-later community instead focus on simpler, foundational programs and concepts such as sequence and control flow.
@@
Unlike CS1 education, however, NLP may often be available as a single elective in a computer science program. Therefore, NLP educators are forced to balance providing practical, deep learning skills to students while also providing solid foundations in classical NLP in a shorter time-frame, perhaps supporting an interleaved approach to NLP curricula.
Authors; Publisher Year Chapter Layout Deep learning focus
Speech and Language Processing; Dan Jurafsky, James H Martin; - 2024 Fundamental Algorithms (Statistical Models, Neural Models, RNNs, LSTMs.) followed by NLP applications and linguistic tasks. Techniques first. classical: Yes.
@@
Natural Language Processing; Pushpak Bhattacharyya, Aditya Joshi; Wiley 2023 Fundamental techniques first. Then neural. Significant focus on NLP problems. Interleave classical and neural both. Three generations.
@@
Practical Natural Language Processing; Sowmya Vajjala, Bodhisattwa Majumder, Anuj Gupta, Harshit Surana; O’Reilly 2020 NLP Primers: Quick introduction to primary NLP concepts. Very task driven: Classification, IE, Chatbots, Applications to domains Cover deep learning Early. Focus on applications and tasks
@@
Table 1: Deep learning focus in some NLP textbooks listed in the reverse-chronological order.
@@ 4 Observations from NLP textbooks and courses
Table 1 presents a summary of some NLP textbooks, in terms of layouts and their deep learning focus. The books were identified using a search on the UNSW library. The focus was identified using the content plan of the book, along with a surface analysis of the chapters. We observe that nearly all textbooks cover classical approaches, primarily in terms of statistical models. We describe two textbooks in particular. The textbook by Jurafsky-Martin  [^3] https://web.stanford.edu/~jurafsky/slp3/ divides its chapters into fundamental algorithms and NLP applications. The first five chapters introduce regular expressions, one-hot vectors and statistical algorithms like support vector machines and logistic regression. In the subsequent chapters, the book covers word embeddings, then begins to combine LSTM with CRFs in the context of POS tagging. Following that, the book covers Transformer, fine-tuning and prompting/prompt learning Transformer. Bhattacharyya and Joshi (2023) use a different approach. They cover fundamental algorithms for representation learning: computational grammar, probabilistic language modeling, and word2vec/LSTM-based representations. The book visualizes NLP as three generations: rule-based, statistical and neural. Following that, the book introduces Transformer. The subsequent chapters alternate between the approaches in the three generations.
Table 2 illustrates how some NLP courses cover classical topics. We only use examples of courses whose course outlines are publicly available on the internet. Similarly, this is not a complete list either. In universities where there is a deep learning-focused course for NLP, topics such as syntactic parsing or statistical classification are covered in substantial detail. Even for courses that cover little classical content, n-gram language modeling is a topic that is included. The table also shows ‘when is Transformer taught’ which may be viewed as a turning point when deep learning-based approaches assume focus in the course. We view Transformer as a transformative technology in NLP, and do not claim that all of deep learning is Transformer.
@@ 5 Lecture Plan
NLP-UNSW: In the NLP-UNSW course which is spread over 10 weeks, we adopt a hybrid methodology where we interleave deep learning approaches with statistical/rule-based approaches. The hybrid methodology is illustrated in Figure 3. In the first week, we introduce NLP via black-boxes such as spacy, nltk, and HuggingFace pipelines. The students are introduced to NLP tasks with demonstrations of the three libraries. This is to help the students develop an understanding of the task. Similarly, if students aim to only ‘use’ NLP in their projects (which may be sufficient for interdisciplinary projects), these libraries are sufficient. The focus here is also to help them understand the input and output of NLP systems, which we believe is the starting point to understanding NLP.
@@
In week 2, we cover representations of text via one-hot vectors and probabilistic language modeling (statistical approach), and word representation learning as in word2vec and GloVe (deep learning approach). This not only allows students to appreciate the value addition of deep learning approaches but also identify situations in which non-deep learning approaches may be sufficient. In week 3, we focus on Transformer architecture: exposing the students to the architectural details, pseudocode, and code implementations. This is followed by Transformer-based models (encoder, decoder, and encoder-decoder models) in week 4. With this background in neural NLP models, we switch gears and focus on one NLP task every week. Every task is covered using the following steps: the linguistic task and associated challenges, classical approaches, deep learning approaches, and recent advances in the area. The first two allow the students to gain a foundational understanding of the problem, followed by the state-of-the-art in deep learning. The recent advances are catered to students who might be interested in future research in NLP without going into too much depth - to ensure that the content is accessible. The NLP tasks we cover are: sentiment analysis (representing sequence classification), named entity recognition (representing token classification), machine translation (representing sequence-to-sequence tasks), summarization (modified sequence-to-sequence tasks) and bias mitigation.
@@
NLP-IITB: NLP-IITB starts with a focus on sequence labeling via HMM-based POS tagging, followed by tree extraction via probabilistic parsing. In both cases, the mathematical details (driven by probability) go hand in hand with the algorithms (explained through the code). Then, machine translation is introduced, which makes way for an introduction to Transformer and large language models. The subsequent weeks discuss sentiment analysis as a specific NLP task and evaluation metrics in NLP. It must be noted that NLP-IITB is the introductory NLP course while there is a deep learning-specific NLP course at IITB which is often a follow-up to the introductory NLP-IITB. This is not the case for NLP-UNSW (being the only NLP-centric course). As a result, the course needs to balance between the two. This comparison shows how educators may make different choices, depending on the length of the teaching period. As a result, the lecture plan of NLP-IITB uses the spectrum shown in Figure 4. classical approaches allow focusing on the linguistic phenomenon which is captured using a model. The data provides the parameters of the model.
@@
Comparison: In both NLP-UNSW and NLP-IITB, however, we observe that the foundational material serves as a good basis to introduce terminology (for example, types of summarization are abstractive and extractive summarization) and tag sets (as in the case of named entity recognition). The classical approaches build an intuitive understanding of building computational models for the specific NLP task. Appreciating the challenges in the classical approaches leads the students to the need for deep learning approaches. Through handwritten examples in the lecture, we highlight the relationships between the mathematics of classical approaches and Transformer-based approaches (for example, alignment model in statistical MT with cross-attention in Transformer).
@@ 6 Coding Assessments
Both NLP-UNSW and NLP-IITB have individual and group coding assessments. NLP-UNSW has one of each, while NLP-IITB has two individual and one group assessment(s).
@@ 6.1 Course Assignment
NLP-UNSW: The first coding assessment in NLP-UNSW is an individual coding assignment completed by students by the end of week 3. The coding assessment is run in a competition-like environment. The students are given a document defining the problem definition, and a template code that also contains stubs for testing. The students are required to add to the template. Closer to the submission, the students are given test cases to evaluate their code. The topic of the individual assignment needed deliberation. Until that point, we have only started discussing Transformer in the lectures. Therefore, we utilize the individual assignment to assess the classical skills of the students. In the case of this offering, the task was to extract skills from job ads, based on a skills ontology. The assignment was designed in two parts: the first part required the students to use the spacy matcher, while the second part required them to use embedding similarities. The former tests them for their ability to use a classical library while the latter requires them to use HuggingFace models. As a result, the assessment covers the students’ ability to ‘use’ NLP models. Marking the assignment includes an automatic marking component with test cases and manual marking by the tutors. Being a rule-based system, the evaluation of test cases is found to return low precision. Therefore, rather than scoring on individual test cases, we mark the students on precision thresholds.
NLP-IITB: NLP-IITB follows a similar strategy although there are multiple individual assignments. The course convener found it useful to give classical assignments on topics such as POS tagging, parsing and so on. An innovative assignment was also tried out, which met with good student feedback. Given a large dataset of names of cities, the task was to identify a preferred suffix in the name of a city. Students reported multiple deep learning/non-deep learning approaches.